OpenStudy (calculusxy):

MEDAL!!! What do I do when for a solution of a parabola I don't get a real number (like 1, 3) instead I get something like:

OpenStudy (calculusxy):

\(\huge \frac{2 \pm 2\sqrt{1} }{3}\)

OpenStudy (anonymous):

\[\sqrt1=??\]

OpenStudy (calculusxy):

fine, we can replace this to be like 3. i am making this up

OpenStudy (anonymous):

oh ok it is a hypothetical question

OpenStudy (calculusxy):

\[\frac{2 \pm 2\sqrt{1} }{3}\]

OpenStudy (anonymous):

you get what you get if the answer is \[\frac{2\pm\sqrt5}{2}\] then those are the two answers

OpenStudy (anonymous):

most of the time you do not get a rational number

OpenStudy (calculusxy):

so if i don't get a rational number, then i will just have to leave it like that right?

OpenStudy (anonymous):

yes

OpenStudy (calculusxy):

ok thank you.

OpenStudy (anonymous):

you have a specific example?

OpenStudy (calculusxy):

another question.

OpenStudy (anonymous):

kk

OpenStudy (calculusxy):

so if i have (hypothetical): \[\frac{ 4 \pm 8\sqrt{2} }{ 4 }\] then i can simplify it as: \[\frac{ 1 \pm \sqrt{3} }{ 1 }\] or just \(1 \pm \sqrt{3}\)

OpenStudy (anonymous):

not quite

OpenStudy (calculusxy):

oh sorry i forgot to put the 2 in front of the square root

OpenStudy (anonymous):

\[\frac{ 4 \pm 8\sqrt{2} }{ 4 }=\frac{4}{4}\pm\frac{8\sqrt2}{4}\]

OpenStudy (anonymous):

so \[1\pm2\sqrt2\] it means you used the quadratic formula when you should have completed the square

OpenStudy (anonymous):

you have the original equation ?

OpenStudy (calculusxy):

\[1 \pm 2\sqrt{2}\] no i just made a hypothetical question because i have a test on this topic (along with others) tmrw

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

good luck with your exam

OpenStudy (calculusxy):

and but if i have: \[\frac{ 2 \pm 6\sqrt{5} }{ 3 }\] then i just leave it as it is?

OpenStudy (anonymous):

yes

OpenStudy (calculusxy):

ok thank you so much!

OpenStudy (anonymous):

i would

OpenStudy (anonymous):

yw

OpenStudy (anaise):

......